# ISLES’24: Improving final infarct prediction in ischemic stroke using multimodal imaging and clinical data

### v1

@@ Abstract
Accurate estimation of core (irreversibly damaged tissue) and penumbra (salvageable tissue) volumes is essential for ischemic stroke treatment decisions. Perfusion CT, the clinical standard, estimates these volumes but is affected by variations in deconvolution algorithms, implementations, and thresholds. Core tissue expands over time, with growth rates influenced by thrombus location, collateral circulation, and inherent patient-specific factors. Understanding this tissue growth is crucial for determining the need to transfer patients to comprehensive stroke centers, predicting the benefits of additional reperfusion attempts during mechanical thrombectomy, and forecasting final clinical outcomes. This work presents the ISLES’24 challenge, which addresses final post-treatment stroke infarct prediction from pre-interventional acute stroke imaging and clinical data. ISLES’24 establishes a unique 360-degree setting where all feasibly accessible clinical data are available for participants, including full CT acute stroke imaging, sub-acute follow-up MRI, and clinical tabular data. The contributions of this work are two-fold: first, we introduce a standardized benchmarking of final stroke infarct segmentation algorithms through the ISLES’24 challenge; second, we provide insights into infarct segmentation using multimodal imaging and clinical data strategies by identifying outperforming methods on a finely curated dataset. The outputs of this challenge are anticipated to enhance clinical decision-making and improve patient outcome predictions. All ISLES’24 materials, including data, performance evaluation scripts, and leading algorithmic strategies, are available to the research community following https://isles-24.grand-challenge.org/.
@@ 1 Introduction
Accurate segmentation of ischemic stroke lesions is crucial for guiding treatment decisions during acute stages (e.g., determining eligibility for thrombectomy), evaluating outcomes, clinical follow-up, and optimizing therapeutic strategies in sub-acute and chronic stages. The Ischemic Stroke Lesion Segmentation (ISLES) Challenge (https://www.isles-challenge.org/) significantly advances stroke image analysis by bringing together experts in neurointervention, radiology, and computer science from top institutions. Biomedical challenges like ISLES are now considered the de facto gold standard for algorithm comparison by the research community and have been organized multiple times, covering diverse pathologies and imaging modalities, thereby further cementing their role in advancing medical imaging technologies [1, 2, 3, 4, 5, 6].
@@
Building on this rich history, ISLES’24 aims to segment the final stroke infarct using pre-interventional acute stroke data. Unlike standard clinical imaging techniques for core estimation, participants have access to the full CT trilogy (non-contrast CT (NCCT), CT angiography (CTA), and perfusion CT (CTP)); follow-up imaging data (DWI and ADC); and clinical tabular data (demographics, clinical history, laboratory results, and stroke outcomes). This creates a unique 360-degree challenge setting where all feasibly accessible clinical data are available for participants. Additionally, ISLES’24 enables the combination and learning from diverse CT techniques with and without contrast, evaluation of brain tissue changes before and after recanalization treatment through follow-up imaging, and inclusion of variables known to impact ischemic tissue status. This comprehensive approach allows participants to refine segmentation models through multimodal data learning strategies. In brief, the main contributions of this work are three-fold:
@@
    The standardized benchmarking of final, post-treatment stroke infarct segmentation algorithms through the ISLES’24 challenge (https://isles-24.grand-challenge.org/).
@@
    Performance evaluation of multimodal imaging and clinical data strategies over a finely curated stroke dataset, including clinical data, acute (NCCT, CTA, CTP), and follow-up (DWI, ADC) stroke imaging.
  3. 3.
    The open access provision of ISLES’24 materials for the research community, including data, performance evaluation scripts, and leading algorithmic strategies.
@@ 2.1 Final infarct prediction using perfusion CT
Perfusion CT is the reference imaging modality for predicting stroke lesion volumes in clinical settings. The workflow involves estimating perfusion maps using established deconvolution approaches and then deriving penumbra and core volumes through thresholding [13]. Although the preferred perfusion map for initial infart (core) estimation varies among software vendors, cerebral blood flow (CBF) and cerebral blood volume maps have been widely adopted. In the last years, various studies propose deep-learning methods for estimating stroke lesions. For instance, de la Rosa et al. [14] propose a deep-learning enhanced SVD-deconvolution method to estimate core tissue. Robben et al. [15] bypass traditional deconvolution by predicting final infarct volumes using native CT Perfusion images and clinical metadata. Their approach combines diverse inputs—CTP, time and treatment parameters, and CTP-derived vascular functions—into a DeepMedic-based model [16], allowing for the simulation of diverse clinical scenarios. Validation of their model over clinical data shows improved estimation of final infarct compared to traditional deconvolution software. Building on this foundation, Amador et al. [17, 18] introduce a deep spatio-temporal CNN to predict final infarct from raw perfusion data. Their model combines a U-Net-like architecture with temporal convolutional networks to efficiently process 4D data. de Vries et al. present PerfU-Net [19], a spatio-temporal model for estimating core tissue from CTP imaging data. PerfU-Net employs a U-Net-based architecture enhanced with an attention module, allowing it to capture dynamic perfusion patterns. Vision transformers have been introduced for infarct segmentation by de Vries et al. [20] and have been later combined with CNN models through hybrid methods in [21]. Gutierrez et al. [22] propose the prediction of follow-up CT images from CTP data using spatio-temporal models, from which the final infarct can be derived from the generated pseudo-CT images. More recently, by leveraging physics-informed neural networks, de Vries et al. [23, 24] show promising results while using a hybrid data-driven image modeling approach.
@@ 2.2 Beyond ASPECTS: Final infarct segmentation using non-contrast CT
Machine and deep learning strategies for infarct segmentation have increasingly been applied to native NCCT due to its broader availability compared to advanced imaging techniques like CTA and CTP. While NCCT reveals infarcted brain tissue as hypodense areas, these changes typically become visible later than those observed with perfusion imaging. Early ischemic changes on NCCT are challenging due to low signal-to-noise ratio, subtle signals, artifacts, and confounding factors [25]. In this context, Srivatsan et al. [26] proposed an image-processing technique for segmenting early ischemic changes by creating a relative NCCT map that highlights subtle hypoattenuation differences between contralateral hemispheres. Qiu et al. [27] trained a standard U-Net model to segment ischemia. Their model, validated over 100 independent scans, demonstrated high volumetric agreement with DWI-derived annotations. Similarly, El-Hariri et al. [28] assessed the performance of the self-adapting nnU-Net model, showcasing high volumetric agreement with ground truth and intraclass correlations over 80%. In 2021, Pan et al. [29] tested a ResNet-based CNN combined with a maximum a posteriori probability model. Later on, Kuang et al. introduced EIS-Net, a multi-task learning model designed to simultaneously segment early infarcts and score the Alberta Stroke Program Early CT Score (ASPECTS) [30]. EIS-Net employs a 3D triplet convolutional neural network that incorporates original and mirrored NCCT images along with an atlas to improve feature extraction. This model excelled in both segmentation and ASPECTS scoring tasks, showing high volumetric agreement with ground-truth annotations and accurate ASPECTS ratings. Wang et al. [31], instead, showed marginal gains of a SwinUNETR model compared to a traditional U-Net method. More recently, Kuang et al. [32] leveraged circular feature interaction and bilateral difference learning in a hybrid CNN-Transformer model to segment ischemia on NCCT.
@@ 2.3 CTA-based final infarct prediction
CTA is a vital imaging modality in stroke management, primarily used to assess cerebral vasculature and identify large vessel occlusions (LVOs). While CTA is traditionally employed for diagnostic purposes, several studies have revealed the potential of CTA-derived features in infarct prediction [33, 34]. Unlike perfusion CT, which provides direct information about blood flow and tissue viability, CTA-based approaches rely on vascular biomarkers such as collateral flow [35, 36, 37] and ASPECTS [33, 38, 36] to estimate the final infarct size.
Building on these findings, Sheth et al.[39] developed an automated deep learning model called DeepSymNet to identify large vessel occlusions (LVO) and the infarct core from CTA source images. Their method demonstrated an ability to determine the infarct core with an AUC of approximately 0.88. Recently, Hokkinen et al.[40] reported that a CTA-based CNN method showed a moderate correlation with final infarct volumes in the late time window for patients with large vessel occlusion who were successfully treated with endovascular therapy. Giancardo et al. [41] further advanced the field by introducing a deep learning method that incorporates a novel weighted gradient-based approach for stroke ischemic core segmentation using image-level labeling. This approach outperforms nnUNet models trained on voxel-level data. More recently, Palsson et al. [42] introduced a U-Net-based convolutional neural network for lesion prediction, augmented with a spatial and channel-wise squeeze-and-excitation block. Their study demonstrates the feasibility of predicting 24-hour follow-up infarcts using acute CTA imaging, achieving results comparable to those of CTP-based algorithms.
@@ 3 Methods
The dataset used in ISLES’24 has been specially prepared for the challenge. It includes multi-scanner and multi-center data derived from large vessel occlusion ischemic stroke cohorts. The dataset includes acute and sub-acute stroke imaging and clinical (tabular) data. Acute imaging data has been acquired at patient admission and includes the diagnostic CT trilogy: NCCT, CTA, and CTP, as well as CTP-derived perfusion maps (namely CBF, cerebral blood volume, mean transit time, and time-to-maximum a.k.a Tmax). Follow-up imaging data was acquired 2 to 9 days later and included DWI and ADC. Moreover, the clinical data contains demographics (age, sex, etc.), patient history, admission scores (e.g., admission NIHSS), and clinical outcomes (e.g., 3-month mRS). In Figure XXX, an example data case is shown.
@@
The dataset is released in raw and preprocessed formats, thus allowing participants to devise algorithms with diverse degrees of freedom. For anonymization purposes, all scans are defaced based on brain and face masks obtained with TotalSegmentator [45], and the clinical tabular data de-identified by adding $\pm$ 5% random noise to the laboratory variables. Preprocessing of the CTP series has been performed using the FDA-approved clinical software icobrain cva [46, 47]. The 4D CTP series are motion-corrected through image co-registration and temporally resampled at 1 frame/second. Afterwards, perfusion maps are generated using a traditional tracer-kinetics deconvolution algorithm. The CTA scans are released after cropping the images around the head, thus removing the thoracic region when available. Preprocessing of the images has been performed by linearly interpolating and registering all the imaging series to the NCCT scans. Except for the MRI scans, where affine transformations are used, all remaining images are registered following rigid transformations. Registration is performed using the Elastix [48] and NiftyReg [49] toolboxes. Furthermore, MRI scans are skull-stripped using HD-BET [50]. Lesion masks are derived from DWI images using the ISLES’22 ensemble algorithm [12]. Quality control and correction of the lesion masks are performed when needed by medical students (TAB, HPM) supervised by two neuroradiologists (JSK, BW) with more than 10 years of experience.
@@
The dataset (N = 250) is split into train (N = 150) and test subsets (N = 100). The train (test) set contains N = 100 (N = 50) scans from the University Hospital of Munich and N = 50 (N = 50) scans from the University Hospital of Zurich. All the train subset is publicly available, while the test subset is hidden from the public, precluding participants from model overfitting strategies. The ISLES’24 data is structured and maintained following the Brain Imaging Data Structure (BIDS) guidelines [51].
@@
Figure 1: Overview of the ISLES’24 task and data modalities. Participants can exploit 1D/2D/3D/4D final infarct learning strategies. NCCT: non-contrast CT; CT: CT angiography; CTP: perfusion CT; CBV: cerebral blood volume; CBF: cerebral blood flow; MTT: mean transit time; Tmax: time-to-maximum; DWI: diffusion weighted imaging.
3.1 Task
@@
The ISLES’24 challenge task is the segmentation of final, post-treatment infarct brain tissue from pre-treatment imaging and clinical data. Participants can choose which image modalities to use and if using them raw or already preprocessed. Multimodal data approaches can be used, as the data involves 1D, 3D, and 4D modalities. An overview of the challenge task and the involved data modalities is depicted in Figure 1.
Algorithms’ evaluation
@@
The final challenge ranking is determined by evaluating each method’s performance on the unseen test set. The test phase of the challenge is blind and single-shot; participants do not have access to the input images, their output results, or the ground truth images. This setup entirely precludes any manual intervention, visual quality assessments, or case-specific overfitting strategies by the teams.
The ranking is performed as in previous ISLES editions [7, 8, 9, 12]. The ranking scheme is derived from the average ranking position obtained from all the individual rankings by case and by metric. First, for each scan and metric, the participants are ranked. Then, for each participant, the average rank across all the metrics for each case is computed, leading to a ’per-case average rank’ for each team. The final ranking is obtained by averaging these ’per-case average ranks’ across the entire dataset.
@@
The ISLES’24 dataset is accessible at at https://isles-24.grand-challenge.org/. To facilitate team participation and metric quantification, two Git repositories have been pre-released. The first repository, https://github.com/ezequieldlrosa/isles24, provides an introduction to the challenge and includes a Python notebook that demonstrates the computation of performance metrics. The second repository (https://github.com/ezequieldlrosa/isles24-docker-template) offers a template Docker algorithm, designed to streamline the creation of algorithmic Dockers. The F1-score and absolute lesion count difference metrics are computed using the panoptica Python library [53]. Connected components are identified using the cc3d [54] Python package. The final challenge ranking is determined using the rankThenAggregate function from the challengeR [55] package implemented in R [56].

### v2

# ISLES’24: Final Infarct Prediction with Multimodal Imaging and Clinical Data. Where Do We Stand?

@@ Abstract
Accurate estimation of brain infarction (i.e., irreversibly damaged tissue) is critical for guiding treatment decisions in acute ischemic stroke. Reliable infarct prediction informs key clinical interventions, including the need for patient transfer to comprehensive stroke centers, the potential benefit of additional reperfusion attempts during mechanical thrombectomy, decisions regarding secondary neuroprotective treatments, and, ultimately, prognosis of clinical outcomes. This work introduces the Ischemic Stroke Lesion Segmentation (ISLES) 2024 challenge, which focuses on the prediction of final infarct volumes from pre-interventional acute stroke imaging and clinical data. ISLES’24 provides a comprehensive, multimodal setting where participants can leverage all clinically and practically available data—including full acute CT imaging, sub-acute follow-up MRI, and structured clinical information—across a train set of $N=150$ cases. On the hidden test set ($N=98$), the top-performing model, a multimodal nnU-Net-based architecture, achieved a Dice score of $0.285\pm 0.213$ and an absolute volume difference of $21.2\pm 37.2$ mL, underlining the significant challenges posed by this task and the need for further advances in multimodal learning. This work makes two primary contributions: first, we establish a standardized, clinically realistic benchmark for post-treatment infarct prediction, enabling systematic evaluation of multimodal algorithmic strategies on a longitudinal stroke dataset; second, we analyze current methodological limitations and outline key research directions to guide the development of next-generation infarct prediction models.
@@ 1 Introduction
Accurate segmentation of ischemic stroke lesions is crucial for guiding treatment decisions during acute stages (e.g., determining eligibility for thrombectomy), evaluating outcomes, clinical follow-up, and optimizing therapeutic strategies in sub-acute and chronic stages. The Ischemic Stroke Lesion Segmentation (ISLES) Challenge (https://www.isles-challenge.org/) significantly advances stroke image analysis by bringing together experts in neurointervention, radiology, and computer science from leading institutions. Biomedical challenges like ISLES are now considered the de facto gold standard for algorithm comparison by the research community and have been organized multiple times, covering diverse pathologies and imaging modalities, thereby further cementing their role in advancing medical imaging technologies [1, 2, 3, 4, 5, 6, 7, 8].
@@
Building on this rich history, ISLES’24 targets the segmentation of the final (post-treatment) stroke infarct, from pre-interventional acute data. The pace at which the hypoperfused tissue converts into irreversibly damaged is widely heterogeneous among patients, brain regions and time, and is driven by diverse tolerance to ischemia and differences in collateral circulation [15]. Anticipating brain tissue evolution in stroke is critical for optimizing personalized treatment decisions, as it helps determine whether (i) mechanical thrombectomy could prevent or limit infarct growth, and whether additional attempts in cases of partial recanalization have added value, (ii) the patient might benefit from additional neuroprotective therapies [16], or (iii) the evolving infarction is likely to cause malignant brain edema, potentially requiring decompressive hemicraniectomy  [17]. Moreover, such predictions can aid in stratifying patient transfers from remote centers by estimating treatment benefit despite delays in transport to thrombectomy-capable stroke units[18]. The prediction of tissue fate may also facilitate the inference of clinical outcomes (e.g., NIHSS or mRS scores) through biomarkers derived from stroke lesions [19, 20], and it may further support resource planning (for example, patients with more severe stroke lesions and predicted outcomes might require longer stays in the stroke unit before being transferred to a general neurological ward or returned to the referring hospital [21]). Given its potential to directly impact acute stroke management and long-term prognosis estimation, accurate prediction of infarct evolution remains a critical and clinically relevant challenge.
Unlike baseline imaging techniques for core estimation, which solely rely on CTP-based algorithms, ISLES’24 offers a unique, 360-degree challenge setting where most feasibly accessible clinical data are available for participants. This includes the full CT imaging suite -non-contrast CT (NCCT), CT angiography (CTA), and perfusion CT (CTP)- as well as follow-up MRI sequences, such as diffusion-weighted imaging (DWI) and apparent diffusion coefficient (ADC) maps. In addition, comprehensive clinical tabular data are provided, comprising demographics, clinical history, laboratory results, neurological scores, and outcome measures. ISLES’24 enables the integration of CT modalities both with and without contrast, supports the longitudinal assessment of brain tissue changes before and after recanalization therapy using follow-up imaging, and incorporates clinically relevant variables known to affect ischemic tissue viability. Altogether, this multimodal and temporally structured challenge design facilitates the development of more robust and generalizable segmentation models through advanced data fusion and learning strategies. In brief, the main contributions of this work are two-fold:
@@
    The establishment of a standardized benchmark for post-treatment stroke infarct segmentation algorithms through the ISLES’24 challenge (https://isles-24.grand-challenge.org/). This includes an extensive algorithmic evaluation conducted on one of the most comprehensive and, to the best of the authors’ knowledge, the first openly available longitudinal stroke dataset.
@@
    The identification of leading solution strategies and the systematic analysis of current algorithmic limitations, along with a discussion of promising future research directions aimed at advancing infarct prediction models.
@@ 2.1 Final infarct prediction using perfusion CT
Perfusion CT is the worldwide accepted imaging modality for predicting stroke lesion volumes in clinical settings. The workflow involves estimating perfusion maps using established deconvolution approaches and then deriving penumbra and core volumes through thresholding [22]. Although the preferred perfusion map for initial infarct (core) estimation varies among software vendors, cerebral blood flow (CBF) and cerebral blood volume maps have been widely installed. Over the past years, various studies have proposed deep learning methods for estimating stroke lesions. For instance, de la Rosa et al. [23] propose a deep learning-enhanced deconvolution method to estimate core tissue. Robben et al. [18] bypass traditional deconvolution by predicting final infarct volumes using native CT Perfusion images and clinical metadata. Their approach combines diverse inputs—CTP, time, and treatment parameters, and CTP-derived vascular functions—into a DeepMedic-based model [24], allowing for the simulation of diverse clinical scenarios. Validation of their model over clinical data shows improved estimation of the final infarct compared to traditional deconvolution software. Building on this foundation, Amador et al. [25, 26] introduce a deep spatio-temporal CNN to predict final infarct from raw perfusion data. Their model combines a U-Net-like architecture with temporal convolutional networks to efficiently process 4D data. de Vries et al. present PerfU-Net [27], a spatio-temporal model for estimating core tissue from CTP imaging data. PerfU-Net employs a U-Net-based architecture enhanced with an attention module, allowing it to capture dynamic perfusion patterns. Vision transformers have been introduced for infarct segmentation by de Vries et al. [28] and have been later combined with CNN models through hybrid methods in [29]. Gutierrez et al. [30] propose the prediction of follow-up CT images from CTP data using spatio-temporal models, from which the final infarct can be derived from the generated pseudo-CT images. More recently, by leveraging physics-informed neural networks, de Vries et al. [31, 32] show promising results while using a hybrid data-driven image modeling approach.
@@ 2.2 Beyond ASPECTS: Final infarct segmentation using non-contrast CT
Machine and deep learning strategies for infarct segmentation have increasingly been applied to native NCCT due to its broader availability compared to advanced imaging techniques like CTA and CTP. While NCCT reveals infarcted brain tissue as hypodense areas, these changes typically become visible later than those observed with perfusion imaging. Early ischemic changes on NCCT are challenging due to low signal-to-noise ratio, subtle signals, artifacts, and confounding factors [33]. In this context, Srivatsan et al. [34] proposed an image-processing technique for segmenting early ischemic changes by creating a relative NCCT map that highlights subtle hypoattenuation differences between contralateral hemispheres. Qiu et al. [35] trained a standard U-Net model to segment ischemia. Their model, validated over 100 independent scans, demonstrated high volumetric agreement with DWI-derived annotations. Similarly, El-Hariri et al. [36] assessed the performance of the self-adapting nnU-Net model, showcasing high volumetric agreement with ground truth and intraclass correlations over 80%. In 2021, Pan et al. [37] tested a ResNet-based CNN combined with a maximum a posteriori probability model. Later on, Kuang et al. introduced EIS-Net [38], a multi-task learning model designed to simultaneously segment early infarcts and score the Alberta Stroke Program Early CT Score (ASPECTS) [39]. EIS-Net employs a 3D triplet convolutional neural network that incorporates original and mirrored NCCT images along with an atlas to improve feature extraction. This model excelled in both segmentation and ASPECTS scoring tasks, showing high volumetric agreement with ground-truth annotations and accurate ASPECTS ratings. Wang et al. [40], instead, showed marginal gains of a SwinUNETR model compared to a traditional U-Net method. More recently, Kuang et al. [41] leveraged circular feature interaction and bilateral difference learning in a hybrid CNN-Transformer model to segment ischemia on NCCT.
@@ 2.3 CTA-based final infarct prediction
CTA is a key imaging modality in stroke management, primarily used to assess cerebral vasculature and identify large vessel occlusions. While CTA is traditionally employed for diagnostic purposes, several studies have revealed the potential of CTA-derived features in infarct prediction [42, 43]. Unlike perfusion CT, which provides direct information about blood flow and tissue viability, CTA-based approaches rely on vascular biomarkers such as collateral flow [44, 45, 46] and ASPECTS [42, 47, 45] to estimate the final infarct size. More recently, multiphase CTA, which consists of one arterial and two venous phase acquisitions, is gaining attention and is increasingly being used for baseline core estimation [48, 49].
Sheth et al.[50] developed an automated deep learning model called DeepSymNet to identify large vessel occlusions and the infarct core from CTA source images. Their method demonstrated an ability to determine the infarct core with an area under the ROC curve of approximately 0.88. Recently, Hokkinen et al.[51] reported that a CTA-based CNN method showed a moderate correlation with final infarct volumes in the late time window for patients with large vessel occlusion who were successfully treated with endovascular therapy. Giancardo et al. [52] further advanced the field by introducing a deep learning method that incorporates a novel weighted gradient-based approach for stroke ischemic core segmentation using image-level labeling. This approach outperforms nnU-net models trained on voxel-level data. More recently, Palsson et al. [53] introduced a U-Net-based convolutional neural network for lesion prediction, augmented with a spatial and channel-wise squeeze-and-excitation block. Their study demonstrates the feasibility of predicting 24-hour follow-up infarcts using acute CTA imaging, achieving results comparable to those of CTP-based algorithms.
@@ 3 Methods
The ISLES’24 dataset includes multi-scanner and multi-center data derived from large vessel occlusion ischemic stroke cohorts. The dataset includes acute and sub-acute stroke imaging and clinical (tabular) data. Acute imaging data have been acquired at patient admission and include the diagnostic CT trilogy: NCCT, CTA, and CTP, as well as CTP-derived perfusion maps (namely CBF, cerebral blood volume (CBV), mean transit time (MTT), and time-to-maximum of the residue function (Tmax)). The follow-up imaging data were acquired 2 to 9 days later and included DWI and ADC. Moreover, the clinical data contains demographics (age, sex, etc.), patient history, admission scores (e.g., admission NIHSS), and clinical outcomes (e.g., 3-month mRS).
@@
The dataset is released in raw and preprocessed formats, thus allowing participants to devise algorithms with diverse degrees of freedom. For anonymization purposes, all scans are defaced based on brain and face masks obtained with TotalSegmentator [56], and the clinical tabular data is de-identified by adding $\pm$ 5% random noise to the laboratory variables. Preprocessing of the CTP series has been performed using the FDA-cleared clinical software icobrain cva [57, 58]. The 4D CTP series are motion-corrected through image co-registration and temporally resampled at 1 frame/second. Afterwards, perfusion maps are generated using a conventional tracer-kinetics deconvolution algorithm. The CTA scans are released after cropping the images around the head, thus removing the thoracic region when available. Preprocessing of the images has been performed by linearly interpolating and registering all the imaging series to the NCCT scans. Except for the MRI scans, where affine transformations are used, all remaining images are registered following rigid transformations. Registration is performed using the Elastix [59] and NiftyReg [60] toolboxes. Furthermore, MRI scans are skull-stripped using HD-BET [61]. Lesion masks are derived from the follow-up MRI using DeepISLES [14]. Quality control and correction of the lesion masks are performed when needed by medical students (TAB, HPM) supervised by two neuroradiologists (JSK, BW) with more than 10 years of experience. Further details about the ISLES’24 data are available in the corresponding data descriptor [62].
@@
The dataset (N = 248) is split into train (N = 150) and test subsets (N = 98). The train (test) set contains N = 100 (N = 50) scans from the University Hospital of Munich and N = 50 (N = 48) scans from the University Hospital of Zurich. The train subset is publicly available, while the test subset is hidden from the public, precluding participants from model overfitting strategies. The ISLES’24 data is structured and maintained following the Brain Imaging Data Structure (BIDS) guidelines [63].
@@
Figure 1: Overview of the ISLES’24 task and data modalities. Participants can exploit 1D/2D/3D/4D final infarct learning strategies. NCCT: non-contrast CT; CT: CT angiography; CTP: perfusion CT; CBV: cerebral blood volume; CBF: cerebral blood flow; MTT: mean transit time; Tmax: time-to-maximum; DWI: diffusion-weighted imaging.
Challenge set-up
Task
@@
The ISLES’24 challenge focuses on the segmentation of final, post-treatment infarcted brain tissue based on pre-treatment imaging and clinical data. Participants are allowed to utilize multiple imaging modalities, provided either in their raw space, in a preprocessed (i.e., co-registered to a common image space) format, or both. Multimodal data approaches can be leveraged, as the dataset spans multiple dimensions—from 1D tabular clinical data to 3D and 4D imaging modalities. An overview of the challenge task and the associated data modalities is provided in Figure 1.
@@
The final challenge ranking is based on the evaluation of each method’s performance on an unseen test set. In this test phase, participants have no access to the input images, their corresponding outputs, or the ground truth annotations. Each algorithm is executed only once on this dataset. This setup entirely precludes any manual intervention, visual quality assessments, or case-specific overfitting strategies by the teams. The ranking is performed as in previous ISLES editions [9, 10, 11, 14] through a ‘rank then aggregate’ strategy, using averaging as the aggregation operation. The ranking scheme is derived from the average ranking position obtained from all the individual rankings by case and by metric. First, for each scan and metric, the participants are ranked. Then, for each participant, the average rank across all the metrics for each case is computed, leading to a ‘per-case average rank’ for each team. The final ranking is obtained by averaging these ‘per-case average ranks’ across the entire dataset.
Baseline models
Two baseline models are incorporated into the challenge framework and evaluated alongside participant submissions by predicting infarct regions on the test set. Both models utilize CTP-derived rCBF perfusion maps, computed after skull-stripping and ventricular structure removal using TotalSegmentator [56, 67]. The first baseline employs the widely used threshold of rCBF $<$ 30%, which has been consistently validated in the literature and adopted in multiple clinical trials, showing strong agreement with infarct volume [68, 69, 70]. The second baseline explores a broader range of rCBF thresholds, from 28% to 46%, selecting the threshold that maximizes the Dice coefficient on the training set. Dice was chosen as the optimization criterion to provide an alternative baseline focused on spatial overlap rather than volumetric agreement.
Statistical analysis
Rankings are computed based on 1000 bootstrap samples. To compare unpaired distributions, we use the Wilcoxon rank-sum test, while paired comparisons of performance metrics across algorithms are assessed using the Wilcoxon signed-rank test, with Holm correction applied for multiple comparisons. Statistical significance is evaluated at the $\alpha=0.05$ level. To evaluate the volumetric agreement between predicted and ground-truth lesion masks, Pearson correlation and Bland-Altman analysis [71] are used.
@@
The ISLES’24 dataset is accessible at xxx. Two Git repositories have been pre-released to facilitate team participation and metric quantification. The first repository, https://github.com/ezequieldlrosa/isles24, provides an introduction to the challenge and includes a Python notebook that demonstrates the computation of performance metrics. The second repository (https://github.com/ezequieldlrosa/isles24-docker-template) offers a template Docker algorithm, designed to streamline the creation of algorithmic Dockers. The F1-score and absolute lesion count difference metrics are computed using the panoptica Python library [66]. Connected components are identified using cc3d [72]. The final challenge ranking is determined using the rankThenAggregate function from the challengeR [73] package implemented in R [74].